# The Minimax Lower Bound of Kernel Stein Discrepancy Estimation

### v1

@@ A.5 Proof of Theorem 2
Next, we present the construction of the adversarial sequence $P_{n}$. Let $\varphi\in\mathcal{C}_{b}(\mathcal{X})$ be as constructed in Lemma B.4, i.e., satisfying $\mathbb{E}_{P_{0}}[\varphi(X)]=0$ and $\varphi\not\equiv 0$. We construct $P_{n}$ as a perturbation of $P_{0}$ taking the form
@@
with $\epsilon_{n}=cn^{-1/2}$, where the precise value of $c>0$ will be specified later. $\varphi\not\equiv 0$ guarantees that $P_{n}\neq P_{0}$; (A.65) implies that $P_{n}\ll P_{0}$ and the corresponding Radon-Nikodym derivative takes the form
@@ Appendix B AUXILIARY RESULTS
[…]
@@
[…]
@@ Appendix C EXTERNAL STATEMENTS
Lemma C.1 (KL divergence of product measures; p. 85; Tsybakov 2009b).

### v2

@@ A.5 Proof of Theorem 2
Next, we present the construction of the adversarial sequence $P_{n}$. Let $\varphi\in\mathcal{C}_{b}(\mathcal{X})$ be as constructed in Lemma B.4, that is, (i) satisfying $\mathbb{E}_{P_{0}}[\varphi(X)]=0$ and (ii) guaranteeing that there exists $A^{\prime}\in\mathcal{B}(\mathcal{X})$ with positive $P_{0}$-measure such that $\varphi(x)\neq 0$ for all $x\in A^{\prime}$. We construct $P_{n}$ as a perturbation of $P_{0}$ taking the form
@@
with $\epsilon_{n}=cn^{-1/2}$, where the precise value of $c>0$ will be specified later; we also note that $P_{n}\neq P_{0}$ by Lemma B.5. (A.65) implies that $P_{n}\ll P_{0}$ and the corresponding Radon-Nikodym derivative takes the form
@@ Appendix B AUXILIARY RESULTS
Let $(\mathcal{X},\tau_{\mathcal{X}})$ be a topological space, $P_{0}\in\mathcal{M}_{1}^{+}(\mathcal{X})$, and $\varphi_{0}\in\mathcal{C}_{b}(\mathcal{X})$ such that there exists no $c\in\mathbb{R}$ such that $\varphi_{0}=c$ holds $P_{0}$-almost surely. Then there exists $\varphi\in\mathcal{C}_{b}(\mathcal{X})$ such that $\mathbb{E}_{P_{0}}[\varphi(X)]=0$ and there exists a set $A\in\mathcal{B}(\mathcal{X})$ with positive $P_{0}$-measure such that $\varphi(x)\neq 0$ for all $x\in A$.
@@
Since $\varphi_{0}\in\mathcal{C}_{b}(\mathcal{X})$, $\varphi_{0}$ is integrable w.r.t. $P_{0}$ and $\mu_{0}\coloneq\mathbb{E}_{P_{0}}[\varphi_{0}(X)]<\infty$. By centering $\varphi_{0}$ as $\varphi\coloneq\varphi_{0}-\mu_{0}$, one has $\mathbb{E}_{P_{0}}[\varphi(X)]=\mathbb{E}_{P_{0}}[\varphi_{0}(X)]-\mu_{0}=\mu_{0}-\mu_{0}=0$. Also, as $\varphi_{0}$ is not constant $P_{0}$-almost surely, the property of $\varphi=0$ $P_{0}$-almost surely does not hold, implying the existence of the stated $A\in\mathcal{B}(\mathcal{X})$. To see that $\varphi\in\mathcal{C}_{b}(\mathcal{X})$, it suffices to note that (i) $\varphi$ is the sum of continuous functions and thus continuous, and (ii) $\sup_{x\in\mathcal{X}}|\varphi(x)|\leq\sup_{x\in\mathcal{X}}|\varphi_{0}(x)|+|\mu_{0}|<\infty$ by the triangle inequality, hence $\varphi$ is also bounded. ∎
Lemma B.5 (Perturbed measures are distinct).
Assume $P_{0}\in\mathcal{M}_{1}^{+}(\mathcal{X})$, let $\varphi\in\mathcal{C}_{b}(\mathcal{X})$ be such that there exists $A^{\prime}\in\mathcal{B}(\mathcal{X})$ with positive $P_{0}$-measure, such that $\varphi(x)\neq 0$ for all $x\in A^{\prime}$, and define the measure $P_{n}$ as $P_{n}(A)=\int_{A}1+\epsilon_{n}\varphi(x)\mathrm{d}P_{0}(x)$, with $\epsilon_{n}>0$, for all $A\in\mathcal{B}(\mathcal{X})$. Then, $P_{0}\neq P_{n}$.
Proof.
We argue by contradiction. Assume that $P_{0}=P_{n}$. Then,
$P_{0}=P_{n}$ $\overset{\text{(a)}}{\implies}1=\frac{\mathrm{d}P_{n}}{\mathrm{d}P_{0}}=1+\epsilon_{n}\varphi\text{ $P_{0}$-almost surely}\overset{\text{(b)}}{\implies}\varphi=0\text{ $P_{0}$-almost surely},$ (B.108)
where (a) uses the definition of the Radon-Nikodym derivative and (b) follows as $\epsilon_{n}>0$. This contradicts the assumption imposed on $\varphi$, concluding the proof. ∎
@@ Appendix C EXTERNAL STATEMENTS
Lemma C.1 (KL divergence of product measures; p. 85; Tsybakov 2009).